Todrick Company is a merchandiser that reported the following information based on 1,000 units sold $270,000 $ 18,000 $ 180,000 $ 9,000 Sales Beginning merchandise inventory Purchases Ending merchandise inventory Fixed selling expense Fixed administrative expense Variable selling expense Variable administrative expense Contribution margin Net operating income $ 10,800 $ 13,500 $ 54,000 $ 16,200 Required: 1. Prepare a contribution format income statement. 2. Prepare a traditional format income statement. 3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate the contribution margin per unit. 6. Which income statement format (traditional format or contribution format) would be more useful to managers in estimating how net operating income will change in responses to changes in unit sales? Complete this question by entering your answers in the tabs below Req 1 Req 2 Req 3 to 5 Req 6 Prepare a contribution format income statement Todrick Company Contribution Format Income Statement Variable expenses 0 Fixed expenses 0

Req 1 Req 2 Req 3 to5 Req 6 Prepare a traditional format income statement. Todrick Company Traditional Income Statement 0 Selling and administrative expenses: 0

3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate the contribution margin per unit. 3. Selling price per unit 4. |Variable cost per unit 5. Contribution margin per unit

Which income statement format (traditional format or contribution format) would be more useful to managers in estimating how net operating income will change in responses to changes in unit sales? OTraditional income statement Contribution income statement
Answer #1

Answer 1:

Todrick Company Contribution Format Income Statement $270,000 Sales Variable expenses: Cost of Goods sold Merchandise $18,000

Calculations:

Variable administrative expense = Sales - contribution - cost of goods sold - variable selling expenses

= 270000 - 54000 - 189000 - 13500 = $13,500

Fixed selling expenses = Contribution - Net operating income - Fixed administrative expense

= 54000 -16200 - 10800

= $27,000

Answer 2:

Todrick Company Traditional Income Statement $270,000 Cost of Goods sold Merchandise Beginning Inventory Add, Purchases Less, Ending Inventory $18,000 $180,000 $9,000 $189,000 Gross Income $81,000 Selling and administrative expenses $27,000 $13,500 Fixed selling expenses Variable selling expenses Fixed administrative expense $10,800 Variable administrative expense $13,500 $64,800 Net Operating Income $16,200

Answer 3, 4 and 5:

3 Selling price per unit 4 Variable cost per unit 5Contribution margin per unit $270 $216 Ş54

Workings:

Selling price per unit = Total sales / Number of units sold = $270,000 / 1,000 = $270

Variable cost per unit = Total variable cost / Number of units sold = $216,000 / 1,000 = $216

Contribution margin per unit = Total contribution / Number of units sold = $54,000 / 1,000 = $54

Answer 6:

Correct answer is:

Contribution income statement

Explanation:

Contribution income statement helps to provides effect of changes in units sales to change in contribution and to change in net operating income and helps in decision making